OpenStudy (anonymous):

B is the midpoint of A and C. If BC=3x2^2, and AC=54, find the value of x.

OpenStudy (anonymous):

x = 3

OpenStudy (anonymous):

Explain how if I substitute x with 3, it would equal 54?

OpenStudy (anonymous):

ac= 54 and bc is mid of ac so bc - 54/2-27 and acc to eq bc=3x^2=27 divide bothsides by 3 (3x^2)/3=27/3 x^2=9 applying square root x=3